OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

parallel: same slope 3=2(5)+b 3=10+b -7=b y=2x-7

OpenStudy (anonymous):

Thank you! @dalia_lam